OpenStudy (anonymous):

Solve by the sustitution method 7x-8y=2 3x+21=y

hero (hero):

Solve each equation for y: y = 3x + 21 y = (2-7x)/-8 Set y = y (by substitution) Substitute each equation in for y: 3x+21 = (2-7x)/-8 Cross Multiply: -8(3x+21) = 1(2-7x) Distribute: -24x - 168 = 2 - 7x Continue Solving For x: -31x = 170 x = -170/31 Substitute x back into original equation to find y

hero (hero):

Did you figure out what y is?

OpenStudy (anonymous):

no